Lou Lewis borrows $15,000 to be repaid over 12 years at 11 percent. Repayment of principal in the first year is: Use Appendix D to calculate the answer.
Answer #1

Assuming payment is made at the end of each year, as nothing is specified in the question

i = 11%

t = 12 yrs

Loan = 15000

Annual loan payments = 15000 * (A/P, 11%,12)

= 15000 * 0.154027

= 2310.41

Interest payment in first payment = 15000 * 0.11 = 1650

Principal payment in first payment = 2310.41 - 1650 = 660.41